Montague Auto Repair specializes in the repair of foreign car transmissions. To encourage business, Montague offers a 6-month warranty on all repairs. The following data are available for 2019:

Transmissions repaired, 2019 4,500
Expected frequency of warranty claims 0.09 per repair
Actual warranty claims, 2019 $110,000
Estimated warranty liability, 1/1/19 $25,000
Estimated cost of each warranty claim $250

Assume that warranty claims are paid in cash.

Required:
1. Compute the warranty expense for 2019.
2. Prepare the entry to record the payment of the 2019 warranty claims.
3. Conceptual Connection: What is the December 31, 2019, balance in the warranty liability account? Why has the balance in the warranty liability account changed from January 1, 2019?
Answer #1

Solution:-

1.Compute the warranty expense for 2019:-

Particulars Amount
Transmissions repaired, 2019 $4,500
Expected frequency of warranty claims $0.09 per repair
Expected Warranty claims

= 4,500 * 0.09

= $405

Estimated cost of each warranty claim $250
Warranty expense for 2019

= 405 * 250

= $101,250

2.Prepare the entry to record the payment of the 2019 warranty claims:-

Particulars Debit Credit
Actual warranty claims, 2019 $110,000
Cash $110,000
( To record the payments)

3.Conceptual Connection: What is the December 31, 2019, balance in the warranty liability account? Why has the balance in the warranty liability account changed from January 1, 2019?

Particulars Amount
Estimated warranty liability, 1/1/19 $25,000
Warranty expense for 2019 $101,250
Actual warranty claims, 2019 $110,000
Balance on, 31/12/2019

= [ 25,000 + 101,250 ] -  110,000

= 126,250 - 110,000

= $16,250